James has two more sisters than brothers. By how much do James' parents' daughters out number their sons? Please, I need this to be solved by tonight. Thank you.
Is the equation like this: x = brothers, x+2 = sisters
1/x + 1/x+2 = 0?

Answer by Edwin McCravy(20086)

Let B = the number of boys (sons) and G = the number of girls (daughters)

Then James has B-1 brothers (he isn't his own brother!) and G sisters.

sisters = brothers + 2

      G = (B-1)+2

      G = B+1

    G-B = 1

So there is one more daughter than sons.

[Note: there is no way to find how many girls or boys,
as there are many possibilities, however there are always
1 more girl than boys. 

Edwin